OpenStudy (anonymous):

Indicate whether the conjecture is true or false. If false, provide a counterexample. Square root of x^2=[x] (P.S. the x after the equal sign is supposed to have two straight lines on either side of it.)

OpenStudy (anonymous):

\[\sqrt{x^2}=|x|\]?

OpenStudy (anonymous):

do you know what |x| means?

OpenStudy (anonymous):

Nope. That's all it gave me :/

OpenStudy (anonymous):

it means 'mod x' or modulus of x - basically the value of |x| is always positive - so for example of x = -3 them mod x = +3

OpenStudy (anonymous):

if x^ 2 = 9 then square root x^2 = 3 or -3

OpenStudy (anonymous):

it is true